Problem: If 68 g of Zn reacts with 27.5 g of HCl. Use the chemical equation to answer the following questions below.


Zn + 2HCl -> ZnC*l_{2} + H_{2}


1. What is the molar mass of Zn?


2. What is the molar mass of HCI?


3. What coefficient is indicated in HCI?


4. Which substance is limiting reactant?


5. Which substance is excess reactant?


6. What is the molar mass of ZnCl_{2} ?


1
Expert's answer
2022-04-11T05:34:04-0400

Solution:

Balanced chemical equation:

Zn + 2HCl → ZnCl2 + H2


1. What is the molar mass of Zn?

Answer (1): The molar mass of Zn is 65.38 g/mol


2. What is the molar mass of HCI?

Answer (2): The molar mass of HCl is 36.458 g/mol


3. What coefficient is indicated in HCI?

Answer (3): The coefficient is 2


4. Which substance is limiting reactant?

5. Which substance is excess reactant?

Solution (4,5):

Calculate the moles of each reactant:

Moles of Zn = (68 g Zn) × (1 mol Zn / 65.38 g Zn) = 1.040 mol Zn

Moles of HCl = (27.5 g HCl) × (1 mol HCl / 36.458 g HCl) = 0.754 mol HCl


According to the chemical equation above:

1 mol of Zn reacts with 2 mol of HCl

Thus, 1.040 mol of Zn reacts with:

(1.040 mol Zn) × (2 mol HCl / 1 mol Zn) = 2.08 mol HCl

However, initially there is 0.754 of HCl (according to the task)

Therefore, HCl acts as limiting reactant and Zn is excess reactant

Answer (4,5):

The limiting reactant is HCl

The excess reactant is Zn


6. What is the molar mass of ZnCl2?

Answer (6): The molar mass of ZnCl2 is 136.286ol